Solve for x:
2log3(x+7+5) = log4*16

I just cannot show the work to get the left side balanced to the right =9.63?

Answer by lwsshak3(11628)


Solve for x:
2log3^(x+7+5)=log4*16
2log3^(x+12)=log64
log3^(x+12)^2=log64
3^(x+12)^2=64
3^(2x+24)=64
(2x+24)log3=log64
2x+24=log64/log3
2x=(log64/log3-24)
x=(log64/log3-24)/2
x≈-10.1072